The force of a hatchdoor of a submarine of area 400cm2 is 45kcak. Calculate the depth of the door below the sea level of the atmospheric pressure is 4atm


1
Expert's answer
2021-03-30T13:47:56-0400

"Area=400 cm^2=0.04m^2"

"F=45 kCal=45 \times4186.80=188406 N"

Pressure at the door, "P_1=\frac{F}{A}=\frac{188406}{0.04}=4710150 \frac{N}{m^2}"

Atmospheric pressure, "P_2=4 atm=405300 \frac{N}{m^2}"

The depth of of the door can be found as below

"P_1+P_2=pgh"

"h=\frac{P_1+P_2}{pg}"

"h=\frac{4710150+405300}{997 \times9.81}=523.022 m"